# Large-scale Pre-training for Grounded Video Caption Generation

### v2

@@ 6.2 Comparison with the state of the art
iGround. The results on our human-annotated iGround test set are shown in Table 2. We compare the results of the proposed GROVE model with our automatic annotation method (described in Section 3) and (still image) GLaMM [28]. The automatic annotation method is a natural fit for a baseline for this task as it performs image-based grounded captioning followed by video-level aggregation without any training. Comparing with [28] aims to assess the benefits of our method in comparison to still-image grounding. Since [28] runs per frame, it does not provide a video-level caption and noun phrases; these differ across frames (see Sec. 3). Thus, we use predictions of [28] for the center frame of each video. We call this the “Center” set-up in Table 2. The “All” set-up considers video-level caption and all frames of the video for bounding box localisation. GROVE, pre-trained on our automatically annotated HowToGround1M dataset (b.), significantly outperforms the [28] baseline (a.) with an improvement of 20 points in CIDEr and 5.9 points in AP50. This demonstrates that temporal context is crucial for this task. Results further improve with finetuning (c., PT+FT). In the all-frame evaluation, the GROVE model pre-trained on the automatically annotated HowToGround1M dataset (e.) provides significant performance improvements over directly obtaining predictions on the test set using the automatic annotation method (d.). This indicates that the GROVE model can correct (smooth out) some of the noise in the automatic annotations during the large-scale training. Finally, we obtain further major performance improvements by fine-tuning GROVE on our manually-annotated iGround dataset, showing that large-scale pre-training with automatic labels followed by fine-tuning on small-scale but highly accurate dataset (PT+FT, c. and g.) is a good recipe for this task.
@@
ActivityNet-Entities and VidSTG. We adapt the GROVE model for the spatio-temporal video grounding task on the ActivityNet-Entities and VidSTG datasets. The details are in the supplementary material. Tables 3 and 4 show the comparison of GROVE with published results on ActivityNet-Entities and VidSTG, respectively. On VidSTG, GROVE achieves the best performance despite not being designed for this task. On ActivityNet-Entities, GROVE significantly outperforms the GVD baseline, demonstrating its effectiveness even with sparse annotations.
Figure 4: Qualitative examples showing predictions of our GROVE model. Please note that GROVE is able to: (i) produce video-level natural language captions describing the main action in the video; (ii) ground multiple objects; and (iii) produce spatio-temporally consistent bounding box predictions. Please note that the second row shows an example of model prediction that is partly incorrect as the blue box, while temporally consistent, does not depict a “yarn”. More qualitative results including video results are in the supp. mat.
@@
Qualitative Results. We show qualitative results of the GROVE model on several example videos in Figures 1 and  4. More qualitative results are in the supplementary material.
@@ 6.3 Effects of pre-training dataset size
Proposed 12.3 31.7 26.9 19.3 22.5
Alt. Stage 1 12.4 36.8 20.1 14.5 20.9
Alt. Stage 2 11.9 28.7 27.8 17.6 21.5
Alt. Stage 3 12.3 31.7 23.2 16.0 20.8
Table 7: Comparison of our automatic annotation approach vs. its alternatives: “Alt. Stage 1” is based on GIT [37], Llama3 [8] and OWLv2 [23]. In “Alt. Stage 2”, we use the full caption instead of Subject-Verb-Object triplets. “Alt. Stage 3” is based on CoTracker3 [14] for tracking the objects of interest. Results are reported on the iGround validation set.
@@ 6.4 Ablation analysis
Model ablations. We perform ablations of the key components of our GROVE model. We report results on the iGround validation set after pre-training with fine-tuning, which achieves the best results. For the ablations, pre-training is performed on a 50k random subset of the HowToGround1M dataset to keep the ablation costs manageable. In Table 7, we ablate the spatio-temporal adapters as well as investigate the importance of training (unfreezing) the bounding box decoder and projection layers (VL and LQ in Fig. 3). To mitigate overfitting, we keep the visual backbones and the LLM frozen. However, we train the embedding and output layers of the LLM to accommodate the modified vocabulary with special tokens.
@@
Ablations of automatic annotation. We replace each stage of our automatic annotation method with an alternative. In Stage 1, we replace the still-image model [28] with an alternative still-image grounded caption generation method. This approach leverages GIT [37] for frame-level captioning, Llama3 [8] for extracting noun phrases from the caption, and OWLv2 [23] for their bounding box localisation within each frame. To ablate Stage 2, we provide the LLM with full captions from Stage 1 instead of extracting Subject-Verb-Object triplets from the caption to assess the impact of additional context. To ablate Stage 3, we incorporate CoTracker3 [14], a SOTA visual point tracking method to provide temporal association of bounding boxes across frames. Using 5 uniformly sampled frames and their bounding box predictions from Stage 1, we track objects in between with CoTracker3 and associate the resulting tracks with noun phrases from the caption. Results are reported in Table 7 where we compare the alternative automatic annotation methods on the iGround validation set. The alternative Stage 1 performs better in captioning due to GIT’s superior performance but performs noticeably worse for grounding. This is because our Stage 1 still-image grounding model [28] is explicitly trained for grounding, unlike GIT, Llama3, and OWLv2, which may underperform due to various factors–such as Llama3 extracting non-groundable noun phrases or OWLv2 missing objects. Compared to our proposed method, the alternative Stage 2 underperforms across all metrics except AP50. This is because our Stage 1 grounding model [28] produces long captions, which include distractor (non-interacted) objects, leading to a small drop in precision (AP50). The alternative Stage 3 underperforms in grounding due to tracker drift caused by abrupt viewpoint changes. Overall, on average (the last column in Table 7) our proposed method achieves the best performance.
@@ 7 Conclusion
In this work, we introduced a large-scale automatic annotation method to generate densely grounded captions in video, which we used to construct the large-scale HowToGround1M video dataset. Building on this, we developed GROVE, a new model for grounded video captioning, which we pre-trained on this large-scale data. To enable rigorous evaluation, we introduced iGround, a dataset with high-quality manual annotations, which also serves as a fine-tuning resource. Our experiments demonstrate the effectiveness of our pre-training strategy, the importance of scaling the pre-training dataset as well as the benefits of fine-tuning on smaller, high-quality data. Our approach not only sets the state of the art on the new manually annotated iGround dataset but also outperforms existing methods on VidSTG and ActivityNet-Entities benchmarks. We believe this work provides a strong foundation for future research in grounded video captioning and highlights the importance of large-scale pre-training combined with smaller-scale but high-quality supervision.
@@ Acknowledgments.
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90140). This research also received the support of projects EXA4MIND and ERC FRONTIER, funded by the European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreements N° 101092944 and N° 101097822. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Furthermore, this work was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI projet), and the ANR project VideoPredict (ANR-21-FAI1- 0002-01). Cordelia Schmid would like to acknowledge the support by the Körber European Science Prize.
Supplementary material
@@ Appendix A Additional qualitative results
Figures 7 and 8 show qualitative results of our GROVE model (Section 4 in the main paper), pre-trained on the HowToGround1M dataset and finetuned on the iGround training set (2000 examples). The results are shown on the iGround test set. In the figures’ captions we discuss some of the benefits of our model. Additional qualitative results showcasing the predictions of our approach overlaid over the input videos are shown in the supplementary video.
Appendix B Dataset Statistics
@@
Figure 7: Qualitative results of our GROVE model on the (unseen) iGround test set. The colour-coded sentence fragments are spatio-temporally localised in the video with the bounding boxes colour coded with the same colour. The results demonstrate that: (i) our model can localise even small objects such as a pen or a tooth brush; (ii) objects are consistently labelled across frames despite changes of viewpoint or scale; (iii) the model focuses on the human and the interacted objects; (iv) the model can successfully ground multiple objects in the video. Additional results are shown in the supplementary video.
@@
Figure 8: Additional qualitative results of our GROVE model on the (unseen) iGround test set. The colour-coded sentence fragments are spatio-temporally localised in the video with the bounding boxes colour coded with the same colour. In addition to the model’s properties discussed in Fig. 7, GROVE is capable of predicting whether an object is present in a certain frame via the temporal objectness head; in the second example there are no bounding box predictions for the hand in the first three frames while in the fourth example there are no predictions for the hand and the screwdriver in the second and fifth frame. Additional results are shown in the supplementary video.
Appendix C Details of the GROVE model
@@
Model architecture. Figure 3 in the main paper shows the different components of our approach. The Global Video Encoder, $\mathcal{V}_{e}(\cdot)$, outputs video features, $o_{e}$, which are pooled spatio-temporally, resulting in the video prompts. These are projected to a language embedding space with $VL(\cdot)$. The LLM, $\mathcal{LM}(\cdot)$, ingests a multimodal prompt consisting of video and language tokens. The LLM is prompted to generate a caption for the video by tagging the noun phrases that correspond to objects and appending them with detection tokens (shown with red and green in the LLM’s generated caption in Figure 3 The LLM’s output hidden states that correspond to the generated caption are projected to queries (using $LQ(\cdot)$). The queries corresponding to the detection tokens are fed to the bounding box decoder $D(\cdot)$. The Grounding Video Encoder, $\mathcal{V}_{g}(\cdot)$, outputs fine-grained video features, which are also fed to the decoder. The decoder performs cross-attention frame-wise between the queries and the outputs of $\mathcal{V}_{g}(\cdot)$, $o_{g}$, which are used as keys/values. Finally, the prediction heads output bounding box predictions and temporal objectness scores for each object at each frame. This objectness score is used to predict the presence/absence of the object in each video frame and is of major importance for the grounded video caption generation task. Details about the visual backbones $\mathcal{V}_{e}(\cdot)$ and $\mathcal{V}_{g}(\cdot)$ as well as details about the LLM $\mathcal{LM}(\cdot)$ including the format of its multimodal inputs and its vocabulary are given next.
@@
Training/inference. We realise the Global Video Encoder $\mathcal{V}_{e}(\cdot)$ with a CLIP-L [27] model with an input of 336$\times$336 and a patch size of 14. The Grounding Video Encoder $\mathcal{V}_{g}(\cdot)$ is instantiated with a SAM [15] encoder and the bounding box decoder $D(\cdot)$ is a SAM-based decoder, the same as in GLaMM [28]. The LLM $\mathcal{LM}(\cdot)$ is a Vicuna-7B model [6]. During training we keep $\mathcal{V}_{e}(\cdot)$, $\mathcal{V}_{g}(\cdot)$ and $\mathcal{LM}(\cdot)$ frozen. $\mathcal{V}_{g}(\cdot)$ originally takes as input 1024$\times$ 1024 images. As this is too large to fit in memory for videos, we instead use 512$\times$512 video spatial resolution, while we interpolate the positional encodings of $\mathcal{V}_{g}(\cdot)$ and fine-tune them. Adapters are 3D spatiotemporal convolutional layers with a kernel of size $3\times 3\times 3$ and a stride of 1. We apply adapters to every 3 layers of $\mathcal{V}_{e}(\cdot)$ and to all global attention layers of $\mathcal{V}_{g}(\cdot)$. The bounding box head $h_{bb}$ is an MLP with two FC layers and a ReLU activation function in between, while the temporal objectness head $h_{tobj}$ is a linear layer. Both prediction heads employ a sigmoid activation function. We apply a threshold of 0.5 to the temporal objectness scores. Both the adapters and the prediction heads are randomly initialised. We use $T=8$ frames for the videos during both training and testing. During training we perform random sparse sampling of frames by splitting the video in 8 segments and randomly drawing a frame from each segment while during testing we pick the centre frame of each segment.
@@
For VidSTG [43] and ActivityNet-Entities [45], we do not use the temporal objectness head. That is because in VidSTG the spatio-temporal tubes are continuous within the segments’ boundaries, while ActivityNet-Entities provides annotations for a single frame per object and in the rest of the frames the objects might still be present but without annotation, and thus should not be modelled as absent. As the task in VidSTG entails predicting the spatio-temporal bounding boxes given a short description, we provide the short descriptions as input to our GROVE model during both training and inference in a teacher-forcing setup.
Appendix D Details of the automatic annotation method

### v3

@@ 6.2 Comparison with the state of the art
iGround. The results on our human-annotated iGround test set are shown in Table 2. We compare the results of the proposed GROVE model with our automatic annotation method (described in Section 3) and (still image) GLaMM [31]. The automatic annotation method is a natural fit for a baseline for this task as it performs image-based grounded captioning followed by video-level aggregation without any training. Comparing with [31] aims to assess the benefits of our method in comparison to still-image grounding. Since [31] runs per frame, it does not provide a video-level caption and noun phrases; these differ across frames (see Sec. 3). Thus, we use predictions of [31] for the center frame of each video. We call this the “Center” set-up in Table 2. The “All” set-up considers video-level caption and all frames of the video for bounding box localisation. GROVE, pre-trained on our automatically annotated HowToGround1M dataset (b.), significantly outperforms the [31] baseline (a.) with an improvement of 20 points in CIDEr and 6.2 points in AP50. This demonstrates that temporal context is crucial for this task. Results further improve with finetuning (c., PT+FT). In the all-frame evaluation, the GROVE model pre-trained on the automatically annotated HowToGround1M dataset (e.) provides significant performance improvements over directly obtaining predictions on the test set using the automatic annotation method (d.). This indicates that the GROVE model can correct (smooth out) some of the noise in the automatic annotations during the large-scale training. We obtain further major performance improvements by fine-tuning GROVE on our manually-annotated iGround dataset, showing that large-scale pre-training with automatic labels followed by fine-tuning on a small-scale but highly accurate dataset (PT+FT, c. and g.) is a good recipe for this task. Finally, the underperformance of GROVE that is directly fine-tuned on iGround without pre-training (FT, f.) provides further evidence of the importance of pre-training, particularly when the fine-tuning dataset is small.
@@
ActivityNet-Entities and VidSTG. We adapt the GROVE model for the spatio-temporal video grounding task on the ActivityNet-Entities and VidSTG datasets. The details are in appendix E. Tables 44 and 5 show the comparison of GROVE with published results on VidSTG and ActivityNet-Entities. On VidSTG declarative sentences (Table 4), GROVE achieves the best performance despite not being designed for this task. On VidSTG interrogative sentences, (Table 4), GROVE outperforms all previous work, both when it is fine-tuned on VidSTG (✓), but also when it is only pre-trained on HowToGround1M pre-training set (✗). On ActivityNet-Entities (Table 5), GROVE significantly outperforms the GVD baseline, demonstrating its effectiveness even with sparse annotations.
Other datasets. In Table 9, we evaluate GROVE on YouCook-Interactions [37] and GroundingYouTube [4], outperforming the previous SOTA by large margins.
@@
Qualitative Results. We show qualitative results of the GROVE model on several example videos in Figures 1 and  4. More qualitative results are in appendix A.
@@ 6.3 Effects of pre-training dataset size
a. Proposed 12.3 31.7 26.9 19.3 22.5
b. Alt. Stage 1 (F) 12.4 36.8 20.1 14.5 20.9
c. Alt. Stage 1 (V) 16.8 23.1 14.6 14.2 17.2
d. Alt. Stage 2 11.9 28.7 27.8 17.6 21.5
e. Alt. Stage 3 12.3 31.7 23.2 16.0 20.8
Table 8: Comparison of our automatic annotation approach (row a.) vs. several alternatives (rows b.-e.) on the iGround validation set. See text for details.
@@ 6.4 Ablation analysis
Model ablations. We ablate the key components of our GROVE model. We report results on the iGround validation set after pre-training with fine-tuning, which achieves the best results. In Table 8, we ablate the spatio-temporal adapters and the training (unfreezing) of the bounding box decoder and projection layers (VL and LQ in Fig. 3). To mitigate overfitting, we keep the visual backbones and LLM frozen, while training the embedding and output layers of the LLM to accommodate the modified vocabulary.
@@
Ablations of automatic annotation. We ablate each stage of our automatic annotation method and show results in Table 8. In “b. Alt. Stage 1 (F)”, we replace our Stage 1 with the GIT [41] image captioner + Llama-3 [9] noun phrase extractor + OWLv2 [24] object detector, which yields slightly crisper captions but weaker grounding because the captioner and detector are not trained jointly; a video-level variant (VideoLlama-3 [45] instead of GIT), row “c. Alt. Stage 1 (V)”, shows the same pattern. Feeding full captions in Stage 2 (row “d. Alt. Stage 2”) instead of using subject-verb-object triplets (row “a. Proposed”) raises AP50: full-captions yield fewer predictions as the LLM trims its output to the most salient objects, reducing recall but improving precision–hence the slightly higher AP50. Adding CoTracker3 in Stage 3 (row “e. Alt. Stage 3”) further degrades grounding due to tracker drift during viewpoint changes. Our proposed automatic annotation method (row “a. Proposed”) scores the best on average; a more detailed analysis is in appendix B.
@@ 7 Conclusion
We introduced a large-scale automatic annotation method to generate densely grounded captions in video, which we used to construct the large-scale HowToGround1M video dataset. We developed GROVE, a new model for grounded video captioning, which we pre-trained on this large-scale data. To enable rigorous evaluation, we introduced iGround, a dataset with high-quality manual annotations, which also serves as a fine-tuning resource. Our experiments demonstrate the effectiveness of our pre-training strategy, the importance of scaling the pre-training dataset as well as the benefits of fine-tuning on smaller, high-quality data. Our approach not only sets the state of the art on the new manually annotated iGround dataset but also outperforms existing methods on the VidSTG, ActivityNet-Entities, GroundingYoutube and YouCook-Interactions benchmarks. We believe this work provides a strong foundation for future research in grounded video captioning.
@@ Acknowledgments.
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90140). This research also received the support of projects EXA4MIND, ELLIOT, CLARA and ERC FRONTIER, funded by the European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreements N° 101092944, N° 101214398, N° 101136607 and N° 101097822. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Furthermore, this work was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI projet), and the ANR project VideoPredict (ANR-21-FAI1-0002-01). Cordelia Schmid would like to acknowledge the support by the Körber European Science Prize.
Appendix
To complement the main paper, this appendix assembles additional results, analyses, and implementation details. Sections A and B provide additional qualitative visualizations and expanded quantitative metrics, respectively. Section C discusses the model and data limitations. Section D reports comprehensive dataset statistics. Section E details the GROVE architecture and training setup, while Section F describes our automatic annotation pipeline. Section G outlines the human-annotation protocol, and Section H lists the exact prompts used to curate spatio-temporally grounded captions.
@@ Appendix A Additional qualitative results
Figures 7 and 8 show qualitative results of our GROVE model (Section 4), pre-trained on the HowToGround1M dataset and finetuned on the iGround training set (2013 examples). The results are shown on the iGround test set. In the figures’ captions we discuss some of the benefits of our model. Additional qualitative results showcasing the predictions of our approach overlaid over the input videos are shown in the supplementary video (available at https://ekazakos.github.io/grounded_video_caption_generation/). Figure 9 shows the main failure modes of our model.
Appendix B Additional quantitative results
Detailed analysis for the ablations of automatic annotation. We replace each stage of our automatic annotation method with an alternative. Results are shown in Table 8. In Stage 1, we replace the still-image model [31] with an alternative still-image grounded caption generation method. This approach leverages GIT [41] for frame-level captioning, Llama3 [9] for extracting noun phrases from the caption, and OWLv2 [24] for their bounding box localisation within each frame. We call this alternative “b. Alt. Stage 1 (F)”. We also evaluate a video-level variant “c. Alt. Stage 1 (V)”, where we replace the GIT captioner with VideoLlama3 [45]. To ablate Stage 2 (“d. Alt. Stage 2”), we provide the LLM with full captions from Stage 1 instead of extracting Subject-Verb-Object triplets from the caption to assess the impact of additional context. To ablate Stage 3 (“e. Alt. Stage 3”), we incorporate CoTracker3 [15], a SOTA visual point tracking method to provide temporal association of bounding boxes across frames. Using 5 uniformly sampled frames and their bounding box predictions from Stage 1, we track objects in between with CoTracker3 and associate the resulting tracks with noun phrases from the caption.
Results are reported in Table 8, where we compare the alternative automatic annotation methods on the iGround validation set. The frame-level alternative Stage 1 (row b.) performs better in captioning due to GIT’s superior performance but performs noticeably worse for grounding. This is because our Stage 1 still-image grounding model [31] is explicitly trained for grounding, unlike GIT, Llama3, and OWLv2, which are not trained jointly and may underperform due to various factors–such as Llama3 extracting non-groundable noun phrases or OWLv2 missing objects. A similar trend is observed for the video-level alternative Stage 1 (row c.). Compared to our proposed method, the alternative Stage 2 (row d.) underperforms across all metrics except AP50. This is because the full-caption input yields fewer predictions as the LLM trims its output to the most salient objects, reducing recall but improving precision–hence the slightly higher AP50. In contrast, the SVO-based input in our proposed automatic annotation method leads to slightly longer captions (12 vs.11 words) with more noun phrases (3.3 vs. 3.0), leading to more object predictions and higher recall. This reflects a typical precision-recall trade-off. The alternative Stage 3 (row e.) underperforms in grounding due to tracker drift caused by abrupt viewpoint changes. Overall, on average (the last column in Table 8), our proposed method achieves the best performance.
Comparison with the state of the art on YouCook-Interactions and GroundgYouTube datasets. In Table 9, we evaluate GROVE on YouCookInteractions and GroundingYouTube datasets, outperforming the previous SOTA by large margins.
Method YouCook-Interactions GroundingYouTube
What When and Where (S3D) [4] 53.98 60.62
What When and Where (CLIP) [4] 58.35 56.98
GROVE 68.67 72.14
Table 9: Comparison with SOTA on YouCook-Interactions [37] and GroundgYouTube[4] datasets.
Method METEOR CIDER AP50 Recall
All Auto. annotation 12.3 31.7 26.9 19.3
GROVE 19.7 92.6 42.0 26.9
Hard Auto. annotation 08.1 07.3 22.3 14.7
GROVE 14.4 41.3 36.0 18.9
Table 10: Results of our automatic annotation method (Auto. annotation) and the complete proposed model (GROVE) on the entire iGround validation set (All) and for a subset (about 10% of data) with challenging similar referring expressions (Hard).
Appendix C Limitations
Although our proposed datasets and model advance the state of the art in grounded video captioning and spatio-temporal sentence grounding, they also reveal avenues for future exploration, stemming from the following limitations.
Scaling to long videos. Despite achieving state-of-the-art results on VidSTG by running inference in a sliding-window manner over videos up to three minutes, the training phase remains memory-bound: we can supply the model with only eight frames per clip. This is sufficient for the short clips in HowToGround1M and iGround (8-10 seconds), where eight frames corresponds to about 1 fps sampling. For VidSTG’s much longer videos, however, uniform sampling of only eight frames introduces large gaps between frames and prevents the model from seeing fine-grained temporal dependencies during training. Closing this discrepancy will require methods that can train directly on larger frame spans or more efficient representations of extended temporal context (e.g. memory).
Complex referring expressions. We examined the iGround validation set and discovered that roughly $10\%$ of its videos contain more than one object whose referring expressions (and appearance) are highly similar. We designate this challenging portion of the data as the “Hard” subset. In Table 10, we compare GROVE with our automatic annotation method in this subset. Although GROVE still surpasses the automatic annotation method on this subset, the marked drop in performance of both methods on the “Hard” subset (comparing to “All”, i.e. the full validation set) reveals that reliably disambiguating closely related referring expressions remains still a challenge. These results suggest opportunities to refine both the model architecture and the automatic annotation method to better handle such fine-grained cases.
Appendix D Dataset Statistics
@@
Figure 7: Qualitative results of our GROVE model on the (unseen) iGround test set. The colour-coded sentence fragments are spatio-temporally localised in the video with the bounding boxes colour coded with the same colour. The results demonstrate that: (i) our model can localise even small objects such as a pen or a tooth brush; (ii) objects are consistently labelled across frames despite changes of viewpoint or scale; (iii) the model focuses on the human and the interacted objects; (iv) the model can successfully ground multiple objects in the video. Additional results are shown in the supplementary video (available at https://ekazakos.github.io/grounded_video_caption_generation/).
@@
Figure 8: Additional qualitative results of our GROVE model on the (unseen) iGround test set. The colour-coded sentence fragments are spatio-temporally localised in the video with the bounding boxes colour coded with the same colour. In addition to the model’s properties discussed in Fig. 7, GROVE is capable of predicting whether an object is present in a certain frame via the temporal objectness head; in the second example there are no bounding box predictions for the hand in the first three frames while in the fourth example there are no predictions for the hand and the screwdriver in the second and fifth frame. Additional results are shown in the supplementary video (available at https://ekazakos.github.io/grounded_video_caption_generation/).
Figure 9: Qualitative results for the main failure modes of our GROVE model on the (unseen) iGround test set. The colour-coded sentence fragments are spatio-temporally localised in the video with the bounding boxes colour coded with the same colour. We identify four main failure modes: (i) temporal objectness mispredicts the presence of an object (first row, last frame for the knife), (ii) inaccurate predictions of object location (second row, third and last frames for the spatula), (iii) misclassification of object (third row, model predicts “radish” for the pumpkin), and (iv) misclassification of action (last example, model predicts “watering” for planting).
Appendix E Details of the GROVE model
@@
Model architecture. Figure 3 shows the different components of our approach. The Global Video Encoder, $\mathcal{V}_{e}(\cdot)$, outputs video features, $o_{e}$, which are pooled spatio-temporally, resulting in the video prompts. These are projected to a language embedding space with $VL(\cdot)$. The LLM, $\mathcal{LM}(\cdot)$, ingests a multimodal prompt consisting of video and language tokens. The LLM is prompted to generate a caption for the video by tagging the noun phrases that correspond to objects and appending them with detection tokens (shown with red and green in the LLM’s generated caption in Figure 3). The LLM’s output hidden states that correspond to the generated caption are projected to queries (using $LQ(\cdot)$). The queries corresponding to the detection tokens are fed to the bounding box decoder $D(\cdot)$. The Grounding Video Encoder, $\mathcal{V}_{g}(\cdot)$, outputs fine-grained video features, which are also fed to the decoder. The decoder performs cross-attention frame-wise between the queries and the outputs of $\mathcal{V}_{g}(\cdot)$, $o_{g}$, which are used as keys/values. Finally, the prediction heads output bounding box predictions and temporal objectness scores for each object at each frame. This objectness score is used to predict the presence/absence of the object in each video frame and is of major importance for the grounded video caption generation task. Details about the visual backbones $\mathcal{V}_{e}(\cdot)$ and $\mathcal{V}_{g}(\cdot)$ as well as details about the LLM $\mathcal{LM}(\cdot)$ including the format of its multimodal inputs and its vocabulary are given next.
@@
Training/inference. We realise the Global Video Encoder $\mathcal{V}_{e}(\cdot)$ with a CLIP-L [30] model with an input of $\times336$ and a patch size of 14. The Grounding Video Encoder $\mathcal{V}_{g}(\cdot)$ is instantiated with a SAM [16] encoder and the bounding box decoder $D(\cdot)$ is a SAM-based decoder, the same as in GLaMM [31]. The LLM $\mathcal{LM}(\cdot)$ is a Vicuna-7B model [7]. During training we keep $\mathcal{V}_{e}(\cdot)$, $\mathcal{V}_{g}(\cdot)$ and $\mathcal{LM}(\cdot)$ frozen. $\mathcal{V}_{g}(\cdot)$ originally takes as input $\times 1024$ images. As this is too large to fit in memory for videos, we instead use $\times512$ video spatial resolution, while we interpolate the positional encodings of $\mathcal{V}_{g}(\cdot)$ and fine-tune them. Adapters are 3D spatiotemporal convolutional layers with a kernel of size $3\times 3\times 3$ and a stride of 1. We apply adapters to every 3 layers of $\mathcal{V}_{e}(\cdot)$ and to all global attention layers of $\mathcal{V}_{g}(\cdot)$. The bounding box head $h_{bb}$ is an MLP with two FC layers and a ReLU activation function in between, while the temporal objectness head $h_{tobj}$ is a linear layer. Both prediction heads employ a sigmoid activation function. We apply a threshold of 0.5 to the temporal objectness scores. Both the adapters and the prediction heads are randomly initialised. We use $T=8$ frames for the videos during both training and testing. During training we perform random sparse sampling of frames by splitting the video in 8 segments and randomly drawing a frame from each segment while during testing we pick the centre frame of each segment.
@@
Details of VidSTG and ActivityNet-Entities experiments.
For VidSTG [48] and ActivityNet-Entities [50], we do not use the temporal objectness head. That is because in VidSTG the spatio-temporal tubes are continuous within the segments’ boundaries, while ActivityNet-Entities provides annotations for a single frame per object and in the rest of the frames the objects might still be present but without annotation, and thus should not be modelled as absent. As the task in VidSTG entails predicting the spatio-temporal bounding boxes given a short description, we provide the short descriptions as input to our GROVE model during both training and inference in a teacher-forcing setup. For evaluating GROVE on VidSTG without observing any VidSTG data during training (GROVE with FT: ✗, Table 4), we pre-train GROVE on HowToGround1M. Each HowToGround1M caption is rewritten–by prompting Llama-3–into both of VidSTG’s sentence styles, declarative and interrogative. Every transformed sentence is then paired with a single bounding box per frame, chosen as the box of the first subject or object it mentions. This supervision reshapes HowToGround1M’s annotation distribution to mirror VidSTG’s, allowing GROVE to achieve strong performance without relying on any VidSTG training data.
Appendix F Details of the automatic annotation method